OpenStudy (anonymous):

Given that alpha is in quadrant II, beta is in quadrant III, sin(alpha)=3/4, and tan(beta)=4/3 find: a) cos(alpha+beta) b) sin(alpha-beta)

OpenStudy (anonymous):

I would use the sum/difference formulas. \[\cos(\alpha+\beta)=\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The other one is: \[\sin(\alpha-\beta)=\sin(\alpha)\cos(\beta)-\sin(\beta)\cos(\alpha)\]

OpenStudy (anonymous):

You also know that tan(beta)=4/3=sin(beta)/cos(beta)

OpenStudy (anonymous):

That should be enough for this problem :)

OpenStudy (anonymous):

oh ok. I didn't think about it that way, I use Tan=opposite/adjacent.

OpenStudy (anonymous):

You could do and solve for beta :) Whichever way would be easier. I believe your way would after you know the sum/difference formulas I provided :P

OpenStudy (anonymous):

how do I do it with alpha though?

OpenStudy (anonymous):

use a^2+b^2=c^2?

OpenStudy (anonymous):

and get the adjacent?

OpenStudy (anonymous):

You can do that. Then you have all the ratios.

OpenStudy (anonymous):

I tried that and got the wrong answer.

OpenStudy (anonymous):

I got that the adjacent is -sqrt(7)

OpenStudy (anonymous):

Negative?

OpenStudy (anonymous):

yeah because its in the second quadrant...

OpenStudy (anonymous):

Ahh, sorry. Yeah.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

no problem :P